B) Hexamethylenediamine (C6H16N2) is one of the starting materials to produce nylon. It can be prepared from adipic acid (C6H10O4) according to the following balanced reaction:

a) What mass of hexamethylenediamine can be produced from 495.5 g of adipic acid?
b) What mass of ammonia is necessary to complete the reaction?
c) If the actual yield of hexamethylenediamine is 301.0 g, determine the % yield for the reaction.
1
Expert's answer
2019-10-07T08:31:58-0400

C6H10O4 + 2NH3 + 2H2 → C6H12N2 + 4H2O

M(C6H10O4) = 116.2 g/mol

M(C6H12N2) = 146.1 g/mol

M(NH3) = 17.0 g/mol

m(C6H10O4) = 495.5 g

mpract(C6H12N2) = 301.0 g


a)     m(C6H12N2)=m(C6H10O4)*M(C6H12N2)/M(C6H10O4)=495.5g*146.1g/mol / 116.2g/mol=623.0g

b)   m(NH3)=m(C6H10O4)*2*M(NH3)/M(C6H10O4)=495.5g*2*17.0g/mol / 116.2g/mol=145.0g

c)    𝜂(C6H12N2)=mpract/mtheor*100%=301.0g/623.0g*100%=48.3%